a convex lens has a focal lenghthof 20 cm in air calculate the change if it immeresed in refractive index 4/3

Grade:

1 Answers

Arun
5 years ago
1/20=(1.5-1)(1/R1+1/R2)
according to second condition
1/f=(1.5/1.3/-1)(1/R1+1/R2)
after dividing both equations u will get answer